# A Picture Is Worth a Graph: Blueprint Debate on Graph for Multimodal Reasoning

### v1

@@ 5.2. Datasets and Metrics
In line with the general setup described in (Zheng et al., 2023b; Lu et al., 2024), we perform our experiments using two extensively adopted multimodal question answering (QA) datasets. These datasets are widely recognized as standard benchmarks, specifically designed to evaluate the performance and effectiveness of models in addressing multimodal reasoning tasks. The two benchmarks are: 1) ScienceQA-IMG (SQA-IMG) (Lu et al., 2022) represents the first multimodal scientific question-answering corpus comprising 21,000 inquiries paired with multiple choices and accompanying images. As a training-free approach, we solely utilize the TEST and DEV partitions of ScienceQA-IMG following prior work (Lu et al., 2022) for comparative assessment. Additionally, the questions within ScienceQA can be categorized into three subdomains: natural sciences (NAT), social sciences (SOC), and linguistic sciences (LAN). 2) MMbench (Liu et al., 2023) offers a more systematic and robust means for reasoning evaluation compared to existing benchmarks such as VQAv2 or COCO Captions. We employ the official data split (MMBench-Dev) and code released by the originating authors. The statistics of the two benchmarks are delineated in Appendix. We report the accuracy metric as determined through a heuristic matching procedure, following the same setting of the official benchmark (Lu et al., 2022).
@@
| GPT-4+CoT (Wei et al., 2022) | - | 71.5 | 75.1 |
| GPT-4+BDoG | - | 77.2 | 79.2 |
| GeminiProVision (Team et al., 2023) | - | 76.5 | 75.2 |
| GeminiProVision+BDoG | - | 81.1 | 81.3 |
5.3. Performance Comparison to SOTA Methods
We evaluate the proposed method by by comparing it against two sets of state-of-the-art approaches as follows:
Open-Source Multimodal LLMs with Relatively Moderate Parameters including MiniGPT-4 (Zhu et al., 2024), Qwen-VL and Qwen-VL- Chat (Bai et al., 2023), CogVLM-Chat (Wang et al., 2023c), mPLUG-Owl2 (Xu et al., 2023), LLaVA-v1.5 (Liu et al., 2024), and InstructBLIP (Dai et al., 2023). These models have parameter scales ranging from 7B to 17B
Closed-Source Multimodal LLMs with Large-Scale Parameters: GPT-3.5 (Wei et al., 2022), GPT-4V (Achiam et al., 2023) and GeminiProVision (Team et al., 2023). Following the general standard, GPT-3.5 and GPT-4 have been incorporated with the CoT (Wei et al., 2022) or DDCoT (Zheng et al., 2023b) (built based on image captioning results). These models are known for their parameter scales above 175B and are considered to have the best performance in most of the literature.
BDoG helps reduce the performance gap between large and small models. It is commonly believed that models with larger parameter scales tend to perform better than smaller ones. This observation generally holds true, as shown in Table 1 for models without BDoG. However, the introduction of BDoG has led to a reduction in the performance gap between these two types of models. This can be seen in the improvement achieved by InstructBLIP, which has experienced a boost of 4.3% and achieves an accuracy of 63.5% on SQA-IMG, comparable to that of GPT-3.5.
BDoG reinforces the multimodal reasoning. Form Table 1, we can also observe the advantage of direct multimodal reasoning (e.g., open-source VL models, and GeminiProVision) over indirect multimodal reasoning (e.g., GPT3.5+CoT and GPT3.5+DDCoT due to their nature of obtaining visual information through image captioning). Even the open-source VL models of the former group achieves comparable performance to those of the latter one, with much smaller parameter scales. With BDoG, which reinforces multimodal reasoning by graph regulation, the performance of direct multimodal reasoning of InstructBLIP and GeminiProVision have been improved by 6.1% and 19.8% on the MMBench dataset.
@@ 5.3. Performance Comparison to SOTA Methods
| InstructBLIP (Dai et al., 2023) | Base | 53.7 | 57.3 | 47.9 | 54.8 | 58.1 | 61.0 | 61.4 | 59.2 | 14.2 | 46.3 | 22.6 | 37.0 | 21.4 | 49.0 | 36.0 |
| + BDoG | 59.7 | 55.6 | 54.2 | 58.1 | 63.1 | 58.2 | 72.7 | 61.4 | 43.7 | 57.8 | 31.5 | 42.7 | 19.5 | 39.1 | 42.6 | |
| + BDoG | 58.1 | 61.3 | 52.1 | 59.0 | 60.6 | 62.6 | 68.2 | 61.5 | 58.8 | 65.5 | 41.2 | 51.2 | 18.6 | 46.1 | 51.1 | |
@@ 5.4. Ablation Study
In order to gain a comprehensive understanding of BDoG, we conduct an ablation study where we decompose BDoG into two variants:
BDoG: we remove the graph regulation and constraints, resulting in a debate-only approach (i.e., vanilla multi-agent debate) for investigating the specific contribution of the debating component of BDoG.
BDoG: we remove the debating rounds, resulting in a graph-based reasoning method for investigating the specific contribution of the graph regulation component of BDoG.
@@
Impact of the debate component: BDoG demonstrates consistent improvements across both benchmarks with a debate-only setting, which encourages LLM agents to collaboratively refine and correct prior responses. For science questions, BDoG facilitates the model’s focus on specific errors, such as direction, size, and position, leading to improved performance in the natural science domain (boosting accuracy from 53.7 to 59.7 for InstructBLIP and 68.9 to 73.3 for GeminiProVision). Additionally, BDoG significantly enhances logical reasoning tasks (improving accuracy from 14.2 to 43.7). However, the debate-only nature has limitations, including trivialization and focus diversion issues. Without the graph regulation, overall performance decreases from 55.8 to 42.6 for InstructBLIP, particularly when addressing questions that require attention to specific attributes and relations.
@@
Impact of the graph regulation: With a graph-regularized knowledge base for the discussion, BDoG also demonstrates consistent improvement of overs the base models on both benchmarks. Compared to the text-based and debate-only method BDoG, it performs evidently better on the MMBench by . This is not a surprise, because MMBench emphasizes more the attribute and relations which challenges a model’s ability on structural reasoning. Although incorporating fact-related graph information proves beneficial in BDoG, the absence of the iteratively refined debate procedure results in decreased performance due to the coarse and distorted extraction of blueprint information.
@@
Impact of combining the debate and graph components: By combining the two components, BDoG achieves gains across nearly all categories. In the ScienceQA-IMG dataset, BDoG exhibits consistent and steady improvements, averaging around 5% compared to the baseline models. This suggests that BDoG is robust and generalizes well for science-related questions. Remarkably, BDoG significantly outperforms the baseline model (InstructBLIP) on the MMBench-Dev set, particularly in the areas of Logical Reasoning (LR) with a margin of 49.1%, Attribute Reasoning (AR) with a margin of 26.6%, and Relation Reasoning (RR) with a margin of 15.2%. BDoG enhances logical reasoning (LR) through a mechanism that refines the reasoning process iteratively, emphasizing the importance of multi-step reasoning rationales. The blueprint graph structure of BDoG, which explicitly models objects, attributes, and relations, contributes to improved reasoning abilities in Attribute Reasoning (AR) and Relation Reasoning (RR). The GeminiProVision model also exhibits comparable performance improvements, with BoG contributing to enhanced fine-grained perception across instances (FP-C), resulting in a gain of 12.2%. This improvement can be attributed to the connections established between various objects within the debate-on-graph framework.
@@
A case study for the iterative improvement on the blueprint: BDoG leverages the advantages of both structured evidence through graph regulation and iterative refinement through debating. This is evident in the consistent improvement observed on the blueprint graph, showcasing the combined benefits of these two components. Figure 3 provides running examples demonstrating the superior reasoning performance of our proposed BDoG framework compared to the BDoG method.
@@
The left case draws from the ScienceQA dataset, testing geographic knowledge and map interpretation. While BDoG correctly answered Dominica is highlighted, it also generated irrelevant information about Dominica’s economic development. This misguided the agents into off-topic discussion, concluding incorrectly with Barbados. In contrast, BDoG concentrated on the question and options, iteratively refining the blueprint entities and relations to arrive at the right answer of Dominica.
@@
The example on the right comes from the MMBench dataset requiring cross-instance perception. As the image contained both candies and jars, it posed a challenge. With BDoG relying on text alone, agreement was rarely reached as responses changed over debate rounds. However, BDoG first generated a blueprint defining image objects and attributes. This established the discussion scope. BDoG then pruned irrelevant candy information, focusing discussion on the specific object - jars. It output the final answer by comparing and connecting the two jar sub-graphs.
In summary, Figure 3 demonstrates that BDoG beats BDoG on both datasets through its blueprint-driven approach. This concentrates graph-based reasoning on salient topics and prunes irrelevant details to arrive at well-supported conclusions.
@@ 5.5. Monitoring The Debating Progress
We evaluated the model’s performance against the termination criteria across multiple debate rounds based on the data in Table 3. Our analysis showed that for models with smaller parameters like InstructBLIP, moving from a single round to two rounds led to significant gains in performance. This improvement was particularly notable when increasing the number of rounds from one to two. However, for larger models that may reach agreement more easily, the performance enhancement was relatively modest when amplifying the number of debate rounds. In general, we found the model’s performance tended to converge within the second or third round. This can be attributed to the underlying reasoning typically being able to answer questions within 2-3 steps.
@@ 5.6. Efficiency Analysis
We further compared the effectiveness versus efficiency of our BDoG framework against BDoG, with results shown in Figure 5. Maintaining concise content focused on key aspects, the graph structure of BDoG demonstrated superior efficiency, requiring approximately 50% less inference time than BDoG. By first generating a blueprint, BDoG defines the scope of the current state, thereby improving model efficiency by filtering irrelevant information. Concurrently, Figure 5 shows BDoG outperformed BDoG in effectiveness, achieving over 5 percentage points higher accuracy than BDoG across three test sets.
This enhanced effectiveness can be attributed to BDoG’s framework concentrating on important knowledge rather than generational textual content without guidance, as in BDoG. Defining a blueprint to establish the discussion boundaries before generating a graph structure enhances both efficiency and effectiveness by restricting the model’s reasoning to salient topics. In summary, Figure 5 illustrates BDoG demonstrated better time performance and predictive power compared to BDoG, validating the utility of our approach in balancing computational cost and predictive accuracy through focused, blueprint-driven debate.

### v2

@@ 5.2. Datasets and Metrics
In line with the general setup described in (Zheng et al., 2023b; Lu et al., 2024), we perform our experiments using two extensively adopted multimodal question answering (QA) datasets. These datasets are widely recognized as standard benchmarks, specifically designed to evaluate the performance and effectiveness of models in addressing multimodal reasoning tasks. The two benchmarks are: 1) ScienceQA-IMG (SQA-IMG) (Lu et al., 2022) represents the first multimodal scientific question-answering corpus comprising 21,000 inquiries paired with multiple choices and accompanying images. As a training-free approach, we solely utilize the TEST and DEV partitions of ScienceQA-IMG following prior work (Lu et al., 2022) for comparative assessment. 2) MMbench (Liu et al., 2023) offers a more systematic and robust means for zero-shot reasoning evaluation compared to existing benchmarks such as VQAv2 (Goyal et al., 2017) or COCO Captions (Chen et al., 2015). We employ the official data split (MMBench-Dev) and code released by the originating authors. We report the accuracy metric through a heuristic matching procedure, following the same setting of the official benchmark (Lu et al., 2022). The statistics of the two benchmarks and detailed settings are delineated in Appendix.
@@
| GPT-4+CoT (Wei et al., 2022) | 175B+ | 71.5 | 75.1 |
| GPT-4+BDoG | 175B+ | 77.2 | 79.2 |
| GeminiProVision (Team et al., 2023) | 175B+ | 76.5 | 75.2 |
| GeminiProVision+BDoG | 175B+ | 81.1 | 81.3 |
@@
| InstructBLIP (Dai et al., 2023) | Base | 53.7 | 57.3 | 47.9 | 54.8 | 58.1 | 61.0 | 61.4 | 59.2 | 19.1 | 54.2 | 34.8 | 47.8 | 24.8 | 56.4 | 44.0 |
| + BDoGDebate | 59.7 | 55.6 | 54.2 | 58.1 | 63.1 | 58.2 | 72.7 | 61.4 | 22.9 | 60.3 | 52.2 | 54.3 | 28.0 | 68.9 | 52.4 | |
| + BDoGGraph | 58.1 | 61.3 | 52.1 | 59.0 | 60.6 | 62.6 | 68.2 | 61.5 | 58.8 | 65.5 | 41.2 | 51.2 | 18.6 | 46.1 | 51.1 | |
@@
5.3. Performance Comparison to SOTA Methods
In contrast to the few-shot methodology, which exhibits susceptibility to the specific examples selected for training, we have opted for the zero-shot setting. This approach circumvents potential biases introduced by a limited sample size, ensuring a more robust and generalizable model. We evaluate the proposed method by by comparing it against two sets of SOTA approaches as follows:
Open-Source Multimodal LLMs with Relatively Moderate Parameters including MiniGPT-4 (Zhu et al., 2024), Qwen-VL and Qwen-VL- Chat (Bai et al., 2023), CogVLM-Chat (Wang et al., 2023c), mPLUG-Owl2 (Xu et al., 2023), LLaVA-v1.5 (Liu et al., 2024), and InstructBLIP (Dai et al., 2023), with parameter scales ranging from 7B to 17B.
Closed-Source Multimodal LLMs with Large-Scale Parameters: GPT-3.5 (Wei et al., 2022), GPT-4V (Achiam et al., 2023) and GeminiProVision (Team et al., 2023). Following the general standard, GPT-3.5 and GPT-4 have been incorporated with the CoT (Wei et al., 2022) or DDCoT (Zheng et al., 2023b) (built based on image captioning results). These models are known for their parameter scales above 175B and are considered to have the best performance.
BDoG helps reduce the performance gap between large and small models. It is commonly believed that models with larger parameter scales tend to perform better than smaller ones. This observation generally holds true, as shown in Table 1 for models without BDoG. However, the introduction of BDoG has led to a reduction in the performance gap between these two types of models. This can be seen in the improvement achieved by InstructBLIP, which has experienced a boost of 4.3% and achieves an accuracy of 63.5% on SQA-IMG, comparable to that of GPT-3.5. Similar results can be found in LLaVA-v1.5 with BDoG which gains the 71.1% accuracy in MMBench, comparable to the GPT-4 model.
BDoG reinforces the multimodal reasoning. Form Table 1, we can also observe the advantage of direct multimodal reasoning (e.g., open-source VL models, and GeminiProVision) over indirect multimodal reasoning (e.g., GPT3.5+CoT and GPT3.5+DDCoT due to their nature of obtaining visual information through image captioning). Even the open-source VL models of the former group achieves comparable performance to those of the latter one, with much smaller parameter scales. With BDoG, which reinforces multimodal reasoning by graph regulation, the performance of direct multimodal reasoning of InstructBLIP and GeminiProVision have been improved by 6.1% and 11.8% on the MMBench dataset.
@@ 5.4. Ablation Study
In order to gain a comprehensive understanding of BDoG, we conduct an ablation study by decomposing BDoG into two variants:
BDoGDebate: we remove the graph regulation and constraints, resulting in a debate-only approach (i.e., vanilla multi-agent debate) for investigating the specific contribution of the debating component of BDoG.
BDoGGraph: we remove the debating rounds, resulting in a graph-based reasoning method for investigating the specific contribution of the graph regulation component of BDoG.
@@
Impact of the debate component: BDoGDebate demonstrates consistent improvements across both benchmarks with a debate-only setting, which encourages LLM agents to collaboratively refine and correct prior responses. For science questions, BDoGDebate facilitates the model’s focus on specific errors, such as direction, size, and position, leading to improved performance in the natural science domain (boosting accuracy from 53.7 to 59.7 for InstructBLIP and 68.9 to 73.3 for GeminiProVision). However, the debate-only nature has limitations, including trivialization and focus diversion issues. Without the graph regulation, overall performance decreases from 55.8 to 52.4 for InstructBLIP, particularly when addressing questions that require attention to multi-hop logistic reasoning (LR) and specific attributes (AR).
@@
Impact of the graph regulation: With a graph-regularized knowledge base for the discussion, BDoGGraph also demonstrates consistent improvement of overs the base models on both benchmarks. Compared to the text-based and debate-only method BDoGDebate, it performs evidently better on the logistic reasoning and attributes reasoning questions by addressing the opinion trivialization and diversion with initialized blueprint. Although incorporating fact-related graph information proves beneficial in BDoGGraph, the absence of the iteratively refined debate procedure results in decreased performance due to the coarse and distorted extraction of blueprint information.
@@
Impact of combining the debate and graph components: By combining the two components, BDoG achieves gains across nearly all categories. In the ScienceQA-IMG dataset, BDoG exhibits consistent and steady improvements, averaging around 5% compared to the baseline models. This suggests that BDoG is robust and generalizes well for science-related questions. Remarkably, BDoG significantly outperforms the baseline model (InstructBLIP) on the MMBench-Dev set, particularly in the areas of Logical Reasoning (LR) with a margin of 44.2%, Attribute Reasoning (AR) with a margin of 17.7%, and Relation Reasoning (RR) with a margin of 3%. BDoG enhances logical reasoning (LR) through a mechanism that refines the reasoning process iteratively, emphasizing the importance of multi-step reasoning rationales. The blueprint graph structure of BDoG, which explicitly models objects, attributes, and relations, contributes to improved reasoning abilities in Attribute Reasoning (AR) and Relation Reasoning (RR). The GeminiProVision model also exhibits comparable performance improvements, with BoG contributing to enhanced fine-grained perception across instances (FP-C), resulting in a gain of 12.2%. This improvement can be attributed to the connections established between various objects within the debate-on-graph framework.
@@
A case study for the iterative improvement on the blueprint: BDoG leverages the advantages of both structured evidence through graph regulation and iterative refinement through debating. This is evident in the consistent improvement observed on the blueprint graph, showcasing the combined benefits of these two components. Figure 3 provides running examples demonstrating the superior reasoning performance of our proposed BDoG framework compared to the BDoGDebate method.
@@
The left case draws from the ScienceQA dataset, testing geographic knowledge and map interpretation. While BDoGDebate correctly answered Dominica is highlighted, it also generated irrelevant information about Dominica’s economic development. This misguided the agents into off-topic discussion, concluding incorrectly with Barbados. In contrast, BDoG concentrated on the question and options, iteratively refining the blueprint entities and relations to arrive at the right answer of Dominica.
@@
The example on the right comes from the MMBench dataset requiring cross-instance perception. As the image contained both candies and jars, it posed a challenge. With BDoGDebate relying on text alone, agreement was rarely reached as responses changed over debate rounds. However, BDoG first generated a blueprint defining image objects and attributes. This established the discussion scope. BDoG then pruned irrelevant candy information, focusing discussion on the specific object - jars. It output the final answer by comparing and connecting the two jar sub-graphs.
@@ 5.5. Monitoring The Debating Progress
We evaluate the model’s performance against the termination criteria across multiple debate rounds based on the data in Table 3. Our analysis shows that for models with smaller parameters like InstructBLIP, moving from a single round to two rounds led to significant gains in performance. This improvement is particularly notable when increasing the number of rounds from one to two. However, for larger models that may reach agreement more easily, the performance enhancement is relatively modest when amplifying the number of debate rounds. In general, we find the model’s performance tend to converge within the second or third round. This can be attributed to the underlying reasoning typically being able to answer questions within 2-3 steps.
@@ 5.6. Efficiency Analysis
We further compare the effectiveness versus efficiency of our BDoG framework against BDoGDebate, as shown in Figure 5. Maintaining concise content focuses on key aspects, the graph structure of BDoG demonstrates superior efficiency, requiring approximately 50% less inference time than BDoGDebate. By first generating a blueprint, BDoG defines the scope of the current state, thereby improving model efficiency by filtering irrelevant information. Concurrently, Figure 5 shows BDoG outperforms BDoGDebate in effectiveness, achieving over 5 percentage higher accuracy than BDoGDebate across three test sets. This enhanced effectiveness can be attributed to BDoG’s concentrating on salient knowledge rather than generational textual content without guidance, as in BDoGDebate.